Training Data

Positive Training (21 cases)
11  13  17  19  2
23  29  3  31  37
41  43  47  5  53
59  61  67  7  71

73

Negative Training (22 cases)
10  12  14  15  16
18  20  21  22  24
25  26  27  28  30
32  33  34  4  6
8  9
Anchors  (5 dimensions)
composite  number  orange prime record

Testing Results

Positive tests  Negative tests
Positive 101, 103, 110
Predictions  107, 109,

79, 83, 89, 91, 97

Negative  36, 38,
Predictions  40, 42,

44, 45, 46, 48, 49

Accuracy  18/19 = 94.74%